# Adaptive Vision-Guided Robotic Arm Control for Precision Pruning in Dynamic Orchard Environments

### v1

@@ 2.4 Evaluation Metrics
The performance of the vision-based controller was evaluated using several quantitative metrics. The trajectory error, $E_{traj}$, is calculated using desired and target joint angles which are then converted into cartesian space and error is computed using the following equation:
$E_{traj}=\sqrt{(x_{c}-x_{t})^{2}+(y_{c}-y_{t})^{2}+(z_{c}-z_{t})^{2}},$ (8)
where $(x_{c},y_{c},z_{c})$ denotes the current end-effector position, and $(x_{t},y_{t},z_{t})$ is the target trajectory location. The success rate, $S$, measures the proportion of trials in which the end-effector successfully aligned with the pruning point with a tolerance of 5 millimeter:
$S=\frac{\text{Number of Successful Trials}}{\text{Total Trials}}\times 100\%.$ (9)
@@ 3 Results and Discussion
Table 1: Summary of Quantitative Results. Key performance metrics showing accuracy, success rate, and response time from simulation trials with the vision-guided controller.
Metric Mean Value Description
End Trajectory Error (mm) 0.23 Error in positioning pruning point
Success Rate (%) 93 Percentage of successful trials
Response Time (s) 2.34 Time to align pruning point
This section presents the experimental results and analysis of the vision-based controller’s performance in apple tree pruning. Table 1 summarizes the results from 20 simulated pruning trials. To test the system’s robustness, the robot was placed in front of the tree trunk, and pruning points were selected at the farthest edges of branches. This setup required the controller to make significant directional adjustments. The system successfully completed 93% of the trials, with an average end trajectory error of 0.23 mm in successful attempts. Figure 3 shows the error trends over time during a typical pruning trial. At the start, the error magnitude decreased rapidly as the proportional controller adjusted the end-effector position. Further refinements using the differential inverse kinematics (IK) approach continued to reduce the error until it reached a steady-state value.
Figure 3: A plot showing the error between the image center and target point converging to zero with successive iterations of the algorithm over time.
The results validate the effectiveness of the proposed vision-based controller in handling precision pruning tasks. The integration of CoTracker3, proportional control, and differential IK enables accurate, reliable, and efficient pruning, even in complex environments. However, the trials also revealed several failure modes that need to be addressed for further improvement. One common failure occurred due to manipulator singularities, where certain joint configurations caused control instability. Some failures also happened when target points were beyond the robot’s kinematic reach. Another limitation was the system’s inability to replan trajectories when occlusions blocked the target, leading to task interruptions. Additionally, in dense branching areas, the tracking algorithm sometimes lost the original target and mistakenly switched to a visually similar branch nearby. These challenges indicate the need for enhancements in trajectory planning and target tracking to improve the system’s overall robustness and reliability in real-world pruning scenarios.
In our real-world dataset, collected from a dormant apple orchard, CoTracker3 demonstrated robust temporal tracking of pruning targets. As illustrated in Figure 4, a close-up view of four consecutive frames highlights how CoTracker3 consistently tracks distinct branch points over time. The model leveraged spatiotemporal attention, ensuring smooth tracking continuity as the camera moves. By incorporating learned feature embeddings, CoTracker3 avoided drift and maintained stable predictions for each tracked keypoint.
(a) Frame at $t_{1}$
(b) Frame at $t_{2}$
(c) Frame at $t_{3}$
(d) Frame at $t_{4}$
Figure 4: Tracking of points over four consecutive frames. Each subfigure represents a frame at time $t_{1},t_{2},t_{3},t_{4}$, showing the consistency of CoTracker3.
Beyond its ability to track points through time, CoTracker3 also provides valuable information into occlusion handling, a common challenge in orchard environments where overlapping branches often obscure pruning targets. As shown in Figure 5, CoTracker3 accurately tracked partially ocluded point. When a point became occluded, the model did not simply discard the tracking instance; rather, it predicted its likely location using information from adjacent frames and spatially correlated features. This ability was helpful to smoothly resume tracking once the point re-emerged into the scene. By using occlusion and continuity information, CoTracker3 provides a reliable mechanism for robotic pruning systems to make informed decisions about when a target is temporarily occluded. This capability is expected to substantially improve the robustness of autonomous pruning operations in dense orchard environments.
@@
Qualitative observation across a set of dormant pruning images showed that the model demonstrated accurate tracking in most cases. However, a few challenging instances were observed. Specifically, the issue was observed when a target completely disappears from few frames due to changes in camera perspective. CoTracker3 occasionally struggled to re-identify previously tracked points when they reappeared. This behavior is attributed to the complexity of orchard environments, where branches often look similar, making re-initialization of lost points difficult. Despite these challenges, CoTracker3 demonstrated a 93% success rate in simulation, confirming its reliability for pruning tasks. For practical deployment, the system required a minimum frame rate of 15-20 Hz to effectively track branch movements. However, the current implementation in simulation operated at lower rates. Future work will focus on optimizing computational efficiency to meet real-time requirements for orchard applications.
4 Conclusion and Future Work
The contemporary agricultural practices rely heavily on manual and intensive human-labor centric processes reducing crop-yield and limiting scalability. Pruning is one of the agricultural techniques that to-date remains a manual process. To this end, this paper proposed an automation algorithm for manual labor-free fruit-tree pruning in dense orchards. The proposed framework utilized CoTracker3, a transformer based robust and real-time feature tracking algorithm, to track the points of interest on fruit-trees and guide a robotic manipulator based pruning tool to the pruning point under extremely challenging and occluded scenes. The framework consists of a feature tracking layer, employing CoTracker3, a planning layer, consisting of a differential inverse kinematics algorithm, and a low-level proportional controller. Extensive experimental and simulation validations demonstrate a success rate of 93% and average trajectory tracking error of 0.23 mm. Future work includes testing the system in actual orchards to assess its performance under varying environmental factors. Furthermore, integrating 3D reconstruction techniques and deformable branch modeling would allow for a more adaptive and precise control strategy, improving the system’s ability to handle dynamic and cluttered environments.

### v2

@@ 2.4 Evaluation Metrics
The performance of the vision-based controller was evaluated using several quantitative metrics across 40 experimental trials. The 3D positioning error, $E_{pos}$, was calculated as the euclidean distance between the current end-effector position and the target pruning point:
$E_{pos}=\sqrt{(x_{\textrm{ee}}-x_{\textrm{target}})^{2}+(y_{\textrm{ee}}-y_{\textrm{target}})^{2}+(z_{\textrm{ee}}-z_{\textrm{target}})^{2}}$ (6)
where $(x_{\textrm{ee}},y_{\textrm{ee}},z_{\textrm{ee}})$ denotes the end-effector position in Cartesian space, and $(x_{\textrm{target}},y_{\textrm{target}},z_{\textrm{target}})$ represents the 3D coordinates of the target pruning point. The mean and standard deviation of end-effector positioning errors are computed to assess overall system accuracy and consistency.
The pixel-space tracking error, $E_{\textrm{pixel}}$, quantifies the magnitude of the error vector in image coordinates:
$E_{\textrm{pixel}}=\sqrt{(x_{c}-x_{t})^{2}+(y_{c}-y_{t})^{2}}$ (7)
where $(x_{c},y_{c})$ is the image center and $(x_{t},y_{t})$ is the tracked point location. The mean and standard deviation of pixel errors provide insight into tracking performance throughout the approach sequence.
Success rates are evaluated at two precision levels: the percentage of trials achieving end-effector positioning within 5 mm of the target pruning point, and the percentage of trials achieving positioning within 10 mm of the target. These metrics provide comprehensive assessment of the system’s precision capabilities for different tolerance requirements in pruning applications.
@@ 3 Results and Discussion
This section presents the experimental results and analysis of the vision-based controller’s performance in orchard tree pruning tasks. The system was evaluated through 40 simulation trials conducted in Gazebo within a simulated orchard environment, with pruning points manually selected across the entire 640×480 pixel image plane at varying positions to assess the controller’s robustness under diverse initial conditions. Figure 3 illustrates the simulation environment showing the camera view with tracked pruning points and the corresponding robot positioning in the Gazebo simulation.
Table 1: Summary of Quantitative Results from 40 Visual Servoing Trials
Metric Value
Total Trials 40
Mean Pixel Error 9.79 px
Std Pixel Error 3.15 px
Mean EE Error 4.28 mm
Std EE error 1.36 mm
Success Rate ($<5$ mm) 77.77%
Success Rate ($<10$ mm) 100%
Figure 3: Simulated camera view of tracked point with statistics displayed. Green colored point indicates tracked pruning point which is to be centered in image frame using UR5e joint-space motion.
Table 1 summarizes the quantitative performance metrics from the complete experimental dataset. The system achieved a mean end-effector positioning error of 4.28 ± 1.36 mm across all trials, demonstrating consistent sub-centimeter accuracy. Visual tracking performance showed a mean pixel error of 9.79 ± 3.15 pixels, indicating effective convergence of the proportional control scheme. Precision analysis reveals that 77.77% of trials achieved positioning accuracy within 5 mm of the target, while 100% of trials achieved accuracy within 10 mm, demonstrating the system’s reliability for precision pruning applications.
Figure 4: Spatial distribution of initial pixel positions colored by final tracking error
CoTracker3 demonstrated robust tracking performance across diverse target locations within the image plane. The algorithm successfully maintained target tracking from initial pixel offsets ranging from 50 to 350 pixels from the image center, with consistent convergence behavior regardless of starting position. Figure 4 illustrates this spatial coverage, showing initial pixel positions color-coded by their final tracking error. The uniform error distribution across the workspace validated CoTracker3’s reliability and the controller’s ability to handle targets positioned anywhere within the camera field of view.
The iterative inverse kinematics solver with damped pseudo-inverse formulation successfully handled all target configurations, achieving 100% success rate within 10 mm tolerance. The adaptive step sizing and joint limit enforcement prevented singularities while maintaining smooth motion profiles throughout the approach sequences. Figure 5 reveals the system’s precision characteristics through the end-effector error distribution. Most trials achieved high precision with errors clustered below 5 mm, while the maximum error reached approximately 10 mm, well within acceptable tolerances for positioning applications. This distribution pattern confirmed the effectiveness of the integrated control approach and demonstrated consistent performance across all tested configurations.
Figure 5: End-effector positioning error distribution with precision thresholds
The proportional gains ($Kp^{x}$, $Kp^{z}$) provided stable convergence without oscillatory behavior across all trials. The depth monitoring automatically halted tracking when the tracked point reached within 20 cm of the camera. The current implementation relies on 2D visual feedback with depth used solely to halt tracking, limiting precision to the achieved 9.79 pixel mean error. The results demonstrate consistent sub-centimeter accuracy across diverse target positions, validating the integrated approach for precision agricultural applications.
4 Conclusion and Future Work
This study presented a vision-based robotic control system for precision tree pruning using CoTracker3 point tracking and iterative inverse kinematics. The system achieved 77.77% accuracy within 5mm tolerance and 100% within 10mm tolerance across 40 simulation trials, with a mean end-effector error of 4.28 ± 1.36 mm, demonstrating the effectiveness of integrating modern point tracking with classical control techniques for agricultural robotics. The instances where the controller failed to achieve sub-5mm precision were primarily attributed to point tracking error when the camera was driven too close to a branch. To evaluate CoTracker3’s robustness beyond simulation, the algorithm was tested on real-world imagery. Figure 6 exemplifies CoTracker3’s occlusion handling capabilities in challenging orchard environments. In frame $t_{1}$, the target pruning point is clearly visible, while in frame $t_{2}$, the same point becomes occluded by overlapping branches. Despite this obstruction, CoTracker3 maintains consistent tracking through its joint attention mechanism, successfully predicting the target’s location using spatiotemporal features from surrounding visible points.
@@ 4 Conclusion and Future Work
While the current simulation trials operated in relatively unoccluded environments, this occlusion-handling capability will become critical in future work involving dense canopies, multiple simultaneous pruning points, and wind-induced branch movement. The demonstrated robustness positions CoTracker3 as an effective foundation for more complex agricultural scenarios where visual obstruction is unavoidable. Future work will address several key limitations to enhance system capabilities. Future plans include integrating end-effector orientation control for angled cuts based on branch geometry, migrating from Gazebo to NVIDIA Isaac Sim for improved physics simulation and realistic force modeling, and incorporating force sensing for reliable contact detection and cut verification. Additionally, implementing collision avoidance algorithms and extending the framework to handle multiple simultaneous pruning points will significantly improve the system’s practical applicability in commercial orchard operations.